A local furniture store is having a terrific sale. They are marking down every price to 45%. If the couch is $440 after the markdown, what was the original price? How much would you save if you bought it at this sale?
Answer by stanbon(75887)

A local furniture store is having a terrific sale. They are marking down every price to 45%. If the couch is $440 after the markdown,
what was the original price?
0.45x = 440
x = 440/0.45 = $977.78
-----------------------------

How much would you save if you bought it at this sale?
Ans: 977.78-440 = $537.78
-----------------
Cheers,
Stan H.
--------------